The cost function for a firm is given by

C(Q)=5+Q2
The firm sells output in a perfectly competitive market, and other firms in the industry sell output at a price of $40.

a) What price should the manager of this firm put on its product?
b) What level of output should be produced to maximize profits?
c) How much profits will be earned?
d) Would you expect the profits to remain in the long run?
1
Expert's answer
2016-01-19T08:44:11-0500
C(Q)=5+Q2, perfectly competitive market, P = $40.

a) As the market is perfectly competitive, all the firms are price takers, so the manager of this firm should put the same price of $40 on its product.
b) To find the level of output, that should be produced to maximize profits, we should find the point, at which MR = MC = P.
MC = C' = 2Q, so 2Q = 40, Q = 20 units.
c) Total profits will be: TP = TR - C = P*Q - 5 - Q^2 = 40*20 - 5 - 20^2 = $395.
d) No, because we will expect the profits to decrease in the long run, until all the firms will earn normal (zero) profits.